Allison:

A cylinder has a radius of 10 m and a height of 8 m. What is the exact volume of the cylinder? 18π m³ 80π m³ 160π m³ 800π m³

Allison:

@Hero

Hero:

@Allison the same formula and steps apply with this one. You start with \(V_c = \pi r^2 h\), input the given values for radius(r) and height(h), then solve for \(V_c\)

Hero:

I'd like to see you try this one on your own. Post all your steps here please.

Allison:

3.14 * 10 * 2 * 8

Hero:

Avoid using 3.14 in this instance. Use π instead.

Allison:

Divided by 3.14 is 160

Hero:

Who told you to divide anything by 3.14? Do you see division symbols in the formula?

Allison:

Triston ._.

Hero:

\(V_c = \pi r^2 h = \pi \times r^2 \times h\)

Allison:

I did that.

Hero:

By the way, \(r^2\) means \(r \times r\) not \(r \times 2\)

Allison:

I KNOW.

Hero:

@Allison, okay but I observe you posted something in the same format as \(r \times 2\) earlier.

Allison:

I got 502.

Hero:

How did you get that? Show me your steps? Remember, you must use \(\pi\) in place of 3.14.

Allison:

3.14 * 10 * 2 * 8

Hero:

The final result will have \(\pi\) in it.

Hero:

You should have \(V_c = \pi \times 10 \times 10 \times 8\)

Allison:

2512 now

Hero:

As I said before \(r^2 = r \times r\) not \(r \times 2\). There's something wrong with your calculator. Do not input \(\pi\) when calculating. Just calculate the \(10 \times 10 \times 8\)

Allison:

180

Hero:

@Allison, what's \(10 \times 10\). Everyone should know how to calculate that WITHOUT a calculator.

Allison:

100!!!!

Hero:

Okay, great, now what's \(100 \times 8\)

Allison:

800.

Hero:

Bingo. Now all you have to do is add the \(\pi\) on the end and remember that the units are cubic.